# Automatic Input Feature Relevance via Spectral Neural Networks

### v1

@@ Abstract
The MNIST dataset As a final test, we applied the spectral features detection method to a reduced version of the well-known MNIST dataset [28]. This latter consists of $28\times 28$ images depicting handwritten digits, ranging from 00 to $9$, to be eventually classified. In our experiments, and to deliver a clear message, we just focused on the subset of images “zeroes” and “ones”. Despite the simplicity of the adopted framework, it is interesting to test the method against a real dataset, where the measure of the relevance of the components is not a priori given. The input vectors of the MNIST dataset has a dimension of $784$, and each component of the vectors represents the intensity of a specific pixel in the corresponding image.
@@
Once again, we trained a spectral feedforward neural network by adjusting the trainable spectral parameters. The hidden structure of the network is identical to that employed for the above applications. Clearly, the input layer is now made of 784 elements, the size of the supplied items. The exit layer is still composed of just two nodes, one associated to the images that display a zero and the other pointing to the ones. As expected, the trained model performs very well in terms of estimated classification accuracy (largest than $99$).
The distribution of the post training eigenvalues associated to the input layer shows a clear peak at zero, signifying that many eigenvalues have been suppressed, thus targeted as ineffective, during training stages. Just a few eigenvalues take non-zero values (see Fig. 7 in Appendix C). This conclusion can be rephrased by stating that many pixels are unnecessary for task resolution.
@@
To figure out how the network exploits the subset of selected pixels for handling the classification task, we can overlay the mask containing only the relevant pixels (the binarized picture displayed as inset in Fig. 3) on individual input (images) data. A few examples of what one gets following this procedure are reported in Fig. (4). Images portraying digit “one” activate a cluster of pixels positioned in the center of the drawing (red circles), as opposed to images with a “zero” displayed, which resonate instead with a complementary set of pixels sorrounding the center. The method is hence capable of identifying a minimal subset of input components which prove crucial for the correct handling of the examined data. By working with a relatively simple dataset, we could elaborate on a sound interpretation, drawn a posteriori, of the process that underlying decision making. This includes to visually isolate the relevant patches of information that are selectively seen by the trained device in operation mode. We anticipate that, for more complex datasets, the posterior interpretation could be less transparent in terms of grasping the qualitative message delivered. Moreover, the method could be also applied as a compression algorithm to store large data in a minimal memory kernel.
@@
Figure 4: Four examples of images from the MNIST dataset as originally presented (top) and after the application of the mask identified by the proposed method (bottom). The red circles identify the central cluster activated by images representing a “one”.
Summing up, and based on the above reported analysis carried out for both mock and real data, we can positively conclude on the adequacy of the proposed features’ detection algorithm. By leveraging on a spectral reparametrization of the optimization protocol, one obtains a set of scalar parameters, the eigenvalues associated to the input nodes of the collection, which provide a sensibe ranking of the input features. The novelty of the method resides in the fact that the obtained ranking is self-consistently and automatically obtained, as an immediate byproduct of the training, without requiring further ad hoc processing of the inspected data. For this reason, and motivated by the positive outcomes of the analysis here reported, we propose the spectral feature detection tool as a viable alternative of other existing schemes appeared in the literature.
Acknowledgements.
This work is supported by #NEXTGENERATIONEU (NGEU) and funded by the Ministry of University and Research (MUR), National Recovery and Resilience Plan (NRRP), project MNESYS (PE0000006) ”A Multiscale integrated approach to the study of the nervous system in health and disease” (DR. 1553 11.10.2022).
Appendix A Spectral Formulation
@@
This Appendix is devoted to provide a few additional details on the spectral approach to the training [20]. As mentioned in the main body of the paper we consider layer $l-1$ and layer $l$ , respectively made of $N$ and $M$ nodes. The state of the system on layer $l-1$ is encapsulated in vector $\textbf{x}^{l-1}=\left(x^{l-1}_{1},\ldots,x^{l-1}_{N}\right)$. The state on layer $l$ is instead stored in vector $\textbf{x}^{l}=\left(x^{l}_{1},\ldots,x^{l}_{M}\right)$. The mapping is implemented as $\textbf{x}^{l}=f(W\textbf{x}^{l-1})$, where $W$ is a $M\times N$ matrix and $f$ The spectral recipe deals with the equivalent transformation $\textbf{z}^{l}=f(A\textbf{z}^{l-1})$, where $\textbf{z}^{l-1}=\textbf{x}^{l-1}=\left(x^{l-1}_{1},x^{l-1}_{2},\ldots x^{l-1}_{N},0,0,\ldots 0\right)$ and the last $M$ elements of vector $\textbf{z}^{l}$ refer to the actual activity displayed on layer $l$. As anticipated in the main body of the paper $A=\Phi\Lambda\Phi^{-1}$ where $\Phi$ and $\Lambda$ are $M\times N$ matrices. The first identifies the eigenvectors of $A$, the latter the associated eigenvalues. More esplicitly we assume:
@@
The weights $w_{ji}$, (associated in the standard formulation to the links of the feedforward network to be trained) are now functions of the spectral attributes of matrix $A$. Following this alternative parametrization, the target of the optimization are the components and the eigenvectors and the corresponding eigenvalues. From optimized eigenvectors and eigenvalues one can compute the weight of the transfer matrix in direct space via equation (7). The markers used for discriminating relevant features are the eigenvalues of $A$ rescaled by a quantity that accounts for the norm of the associated eigenvectors, in accordance with what was done in [26, 25]. In formulae, $\tilde{\lambda}_{i}=\lambda_{i}\left(\sum_{j}\phi_{ji}\right)^{1/2}$. All along the paper we drop the tilde and refer to $\tilde{\lambda}_{i}$ as to the $i$-th eigenvalue of matrix $A$.
Appendix B Training procedure
In this Appendix, we provide detailed information on the training procedures. The optimization process for all three datasets discussed in the main text was conducted using the Adam optimizer [29]. The learning rates were set to 0.01 for the first two datasets and 0.0001 for MNIST. Sparse Categorical Crossentropy was employed as the loss function for all three tasks. The number of epochs was set to 1000 for the Gaussian dataset, 500 for the correlated Gaussian dataset, and 5000 for MNIST. The results presented in the manuscript demonstrated robustness to variations in these hyperparameters. Therefore, the hyperparameters were manually optimized from a restricted set of possible candidates by assessing the model’s performance on a validation set ($10$ of the training set).
Appendix C Datasets
In this Appendix, we clarify the structure of the datasets that we have generated to validate the method. Fig. 5 shows the distributions of the input components for the dependent Gaussian dataset. In the first two panels of Fig. 6 The distributions of the values of the fourth and fifth components of the input vectors in the correlated Gaussian dataset are reported, while in the bottom panel, the same components are plotted against each other.
Figure 5: Distribution of the values of the 20 components of the input vectors in the independent Gaussians dataset. The distributions for the data belonging to the first (resp. second) class are represented in red (resp. blue). For some components, the two classes are indistinguishable, while in some other cases the two distributions are only partially overlapping.
Figure 6: The distributions of the values of the fourth and fifth components of the input vectors in the correlated Gaussian dataset are shown in the first two panels. In the bottom panel, the fifth component is plotted against the corresponding fourth component for some input vectors of the same dataset. Individually, both components contain no information on the vector classes. However, by observing their correlations, a certain degree of asymmetry can be appreciated for data belonging to different classes. By analyzing the two components simultaneously one can therefore acquire useful information to solve the classification task.
Appendix D Post-training eigenvalues distribution for the MNIST dataset
We report here the distribution of the eigenvalues obtained at the end of the training for the MNIST dataset. Specifically, the distribution is shown in panel (a) of Fig. 7 and exhibits a peak at zero, as already mentioned in the main body of the paper. The presence of this peak demonstrates the irrelevance of a large fraction of the input features. Panel (b) of the image 7 shows the $28\times 28$ pixels, the canvas on which images are painted, colored according to the associated eigenvalue score: blue pixels refer to smaller eigenvalues, while yellow pixels stand for large positive values.
Figure 7: In panel (a) the distribution of the first $784$ post-training eigenvalues (normalized to their maximum) for the MNIST datast is reported. Panel (b) shows the input space for the same dataset, where each component (i.e., pixel) is colored according to its relative post-training eigenvalue.

### v2

@@ 3.3 The MNIST dataset
As an additional test, we applied the spectral features detection method to a reduced version of the well-known MNIST dataset [29]. This latter consists of $28\times 28$ images depicting handwritten digits, ranging from 00 to $9$, to be eventually classified. In our experiments, and to deliver a clear message, we first focused on the subset of images “zeroes" and “ones". Despite the simplicity of the adopted framework, it is interesting to test the method against a real dataset, where the measure of the relevance of the components is not a priori given. The input vectors of the MNIST dataset has a dimension of $784$, and each component of the vectors represents the intensity of a specific pixel in the corresponding image.
@@
Figure 5: In panel (a) the distribution of the first $784$ post-training eigenvalues (normalized to their maximum) is reported. Panel (b) shows the input space for the same dataset, where each component (i.e., pixel) is colored according to its relative post-training eigenvalue.
Once again, we trained a spectral feedforward neural network by adjusting the trainable spectral parameters. The hidden structure of the network is identical to that employed for the above applications. Clearly, the input layer is now made of 784 elements, the size of the supplied items. The exit layer is still composed of just two nodes, one associated to the images that display a zero and the other pointing to the ones. As expected, the trained model performs very well in terms of estimated classification accuracy (larger than $99$).
The distribution of the post training eigenvalues associated to the input layer is reported in panel (a) of Fig. 5. A clear peak at zero appears, signifying that many eigenvalues have been suppressed, thus targeted as ineffective, during training stages. Just a few eigenvalues are non-zero. This conclusion can be rephrased by stating that many pixels are unnecessary for task resolution. Panel (b) of the image 5 shows the $28\times 28$ pixels, the canvas on which images are painted, colored according to the associated eigenvalue score: blue pixels refer to smaller eigenvalues, while yellow pixels stand for large positive values.
@@
To figure out how the network exploits the subset of selected pixels for handling the classification task, we can overlay the mask containing only the relevant pixels (the binarized picture displayed as inset in Fig. 6) on individual input (images) data. A few examples of what one gets following this procedure are reported in Fig. (7). Images portraying digit “one" activate a cluster of pixels positioned in the center of the drawing (red circles), as opposed to images with a “zero" displayed, which resonate instead with a complementary set of pixels surrounding the center. The method is hence capable of identifying a minimal subset of input components which prove crucial for the correct handling of the examined data.
@@
Figure 7: Four examples of images from the MNIST dataset as originally presented (top) and after the application of the mask identified by the proposed method (bottom). The red circles identify the central cluster activated by images representing a “one".
To further elaborate on the performance of the proposed method for input feature detection, we repeat the above analysis for the subset of images relative to numbers “five" and and “nine". The results are in line with those reported above, as it can be readily appreciated by visual inspection of Figure 8. With just 14 pixels ($\simeq 1.8\%$ of the total) the network achieves approximately $90\%$ classification accuracy. These pixels are localized in the upper right portion of the image where the "nines" feature a closed circular loop, which is instead lacking in the "fives". This is hence the macroscopic distinctive attribute that is learned by the neural network to deal with the assigned task.
Figure 8: Results for the "five" vs. "nine" case. The neural network was trained to distinguish between images representing the number five and those representing the number nine. Panel (a) displays the distribution of eigenvalues at the end of the training process. A significant peak at zero is observed, indicating that many pixels are not employed for solving the task. Panel (b) illustrates the accuracy trend as a function of the number of pixels provided to the network as input, as follows the reported spectral ranking. Panel (c) presents three example inputs along with the pixels activated after applying the minimal mask. Panel (d) highlights a specific input case where the pixels belonging to the minimal mask and activated for the considered input are marked in red.
Finally, we also report in Figure 9 the recorded accuracy for the complete collection of digits (ranging from 00 to $9$) which compose the entire MNIST dataset. Also in this case, and in analogy with the results obtained for the examined settings in which pair of digits are directly compared, the accuracy raises as a function of the number of input pixels included in the analysis, and ranked according to their spectral importance.
Figure 9: Eigenvalue distribution after training for the full MNIST case (left) and accuracy of the model as the number of pixels taken as input grows (right). The order in which the pixels are inserted follows the absolute value (magnitude) of their corresponding post-training eigenvalues
3.4 Stellar spectra and metallicity prediction
In this Section we report on the performance of the algorithm against a dataset of interest for astrophysics. More specifically,we consider the MILES stellar library [30], composed of 985 stars. The dataset is composed of stellar spectra, namely flux as a function of wavelength, in the range 3525-7500 AA. For each spectrum we consider the associated stellar parameters (effective temperature, gravity, and the abundance of heavy elements known as metallicity) from [31]. A few representative spectra are shown in Figure 10. Because of its spectral type coverage and flux calibration accuracy this stellar library has been extensively used as a basis for constructing stellar population models to analyze galaxy spectra [32, 33] and therefore deriving metallicity and star-formation histories of galaxies.
Figure 10: Plot of some of the stellar spectra composing the MILES dataset [30].
To estimate the metallicity, we here train a spectral feedforward neural network, which takes as an input the displayed intensity of the recorded spectra at various wavelengths. The network architecture consists of three hidden layers comprising 200, 50, and 50 neurons, respectively. As in the spirit of the above, the spectral ranking of the supplied input can guide through the identification of a subset of features deemed relevant for an adequate assessment of the star metallicity.
Figure 11: In the upper-left panel the normalized distribution of the eigenvalues of the input layer for the metallicity prediction task is reported in vertical log scale. Upper-right panel shows the mean absolute error of the model versus the fraction of spectral lines provided as an input. The order in which the lines are inserted follows the absolute value (magnitude) of their corresponding post-training eigenvalues. The bottom panel shows the wavelength ranges in the stellar spectra which are deemed relevant for the metallicity prediction task (grey shadows) along with the first layer eigenvalues (see also Method section)
We chose metallicity as the target because it can be inferred from well-defined absorption lines of metal species. Such lines are, however, generally blended at the resolution of our dataset and in galaxy spectra. A set of “indices”, or wavelength regions, have therefore been defined [34, 35], which are sensitive to various combinations of metallicity, temperature, and gravity. We will compare the feature importances obtained from the machine learning analysis with these classical indices.
[…]
The upper-left panel in Figure (11) shows the normalized eigenvalues distribution of the input layer eigenvalues. As for the other examples discussed above, a clear peak is displayed close to zero (note the logarithmic scale on the $y$-axis) thus implying that several input features play no role for the handled task. In particular, if we compute the prediction error on the test set as a function of the number of spectral lines used as input — ordered according to their corresponding eigenvalues — we observe that the minimum error is reached using only around $30$ lines (see the upper-right panel in Figure 11). In other words, the network is able to make accurate predictions by relying on just $0.7\%$ of the input components. The eigenvalues allow hence to identify a subset of relevant input elements that drive decision making across the network. In this respect, the algorithm is thus successfully tested against a real dataset with non trivial physical implications.
4 Discussion
We have here elaborated on a procedure for feature relevance extraction, which is founded on solid mathematical grounds. The interest of the proposed method resides in that the user is not asked to decide a priori on what it is relevant or not. The assessment on the actual relevance of the provided input feature follows the evaluation of a scalar indicator (one for each of the features) that self consistently gauges the impact of the supplied information in the decision making problem. These parameters are mathematically traced back to the eigenvalues of the transfer operator linking the first two layers. The method is hence capable of self-adjusting to the specific problem under inspection, with no need for ad hoc (possibly biased) input from external operators.
Based on the above reported analysis carried out for both mock and real data, we can positively conclude on the adequacy of the proposed features’ detection algorithm.The eigenvalues associated to the input nodes of the collection provide indeed a sensible ranking of the input features.
By working with a relatively simple dataset, we could elaborate on a sound interpretation, drawn a posteriori, of the process that underlying decision making. This includes to visually isolate the relevant patches of information that are selectively seen by the trained device in operation mode. For more complex datasets, the posterior interpretation could be less transparent in terms of grasping the qualitative message delivered. As an obvious corollary of the analysis, the method could be also applied as a compression algorithm to store large data in a minimal memory kernel.
5 Methods
@@
Spectral Formulation We provide here a few additional details on the spectral approach to the training [22]. As mentioned in the main body of the paper we consider layer $l-1$ and layer $l$ , respectively made of $N$ and $M$ nodes. The state of the system on layer $l-1$ is encapsulated in vector $\textbf{x}^{l-1}=\left(x^{l-1}_{1},\ldots,x^{l-1}_{N}\right)$. The state on layer $l$ is instead stored in vector $\textbf{x}^{l}=\left(x^{l}_{1},\ldots,x^{l}_{M}\right)$. The mapping is implemented as $\textbf{x}^{l}=f(W\textbf{x}^{l-1})$, where $W$ is a $M\times N$ matrix and $f$ The spectral recipe deals with the equivalent transformation $\textbf{z}^{l}=f(A\textbf{z}^{l-1})$, where $\textbf{z}^{l-1}=\textbf{x}^{l-1}=\left(x^{l-1}_{1},x^{l-1}_{2},\ldots x^{l-1}_{N},0,0,\ldots 0\right)$ and the last $M$ elements of vector $\textbf{z}^{l}$ refer to the actual activity displayed on layer $l$. An appropriate projection matrix should be formally considered at the end of each mapping between consecutive layers, which effectively implements the transfer of the activations from bottom to upper elements (while filling the bottom portion of the vector with zeros) of the next vector to be processed.
As anticipated in the main body of the paper $A=\Phi\Lambda\Phi^{-1}$ where $\Phi$ and $\Lambda$ are $M\times N$ matrices. The first identifies the eigenvectors of $A$, the latter the associated eigenvalues. More explicitly we assume:
@@
The weights $w_{ij}$, (associated in the standard formulation to the links of the feedforward network to be trained) are now functions of the spectral attributes of matrix $A$. Following this alternative parameterization, the target of the optimization are the components and the eigenvectors and the corresponding eigenvalues. From optimized eigenvectors and eigenvalues one can compute the weight of the transfer matrix in direct space via equation (7). The markers used for discriminating relevant features are the eigenvalues of $A$ rescaled by a quantity that accounts for the norm of the associated eigenvectors. In formulae, $\tilde{\lambda}_{i}=\lambda_{i}\left(\sum_{j}\phi_{ji}\right)^{1/2}$. All along the paper we drop the tilde and refer to $\tilde{\lambda}_{i}$ as to the $i$-th eigenvalue of matrix $A$.
Training procedure The optimization process for all three datasets discussed in the main text was conducted using the Adam optimizer [38]. The learning rates were set to 0.01 for the first two datasets, 0.0001 for MNIST and 0.001 for the stellar dataset. Sparse Categorical Crossentropy was employed as the loss function for the three classification tasks, while the mean absolute error was used as loss function for the metallicity prediction task. For all tasks under consideration, we applied regularization to both the eigenvalues and eigenvectors of the first layer. The regularization coefficients were set to $10^{-2}$ for the eigenvalue term and $10^{-4}$for the eigenvector term. The number of epochs was set to 1000 for the Gaussian dataset, 500 for the correlated Gaussian dataset, 5000 for MNIST and 1000 for the stellar spectra dataset. Using a large number of training epochs allows us to significantly reduce the regularization term associated with the eigenvalues in the loss function. This, in turn, enables the model to converge to a solution that relies on a minimal subset of input features while still correctly classifying the data. Such sparsity in the input improves the interpretability of the model. It is important to note that employing a large number of epochs is not a stringent requirement: even when operating with fewer training epochs or smaller regularization terms, the eigenvalues of the input layer remain valid indicators of node relevance. However, under these conditions, the model tends to distribute importance across a larger number of input components. While this still yields accurate classification, the interpretation of the internal representations becomes more diffuse and less localized.
As empirical evidence of the above, we report the results for the MNIST dataset and the stellar dataset as obtained for a smaller number of training epochs and reduced regularization factors. By significantly reducing the number of epochs during training on the MNIST dataset, we find a bimodal distribution for the eigenvalues that refer to the input layer with a peak close to zero that is less pronounced as compared to that reported in Sec. (3.3), and a prominent bulk in the region of larger values (See left panel of Fig. 12). This indicates that a greater number of input components are being used to solve the task. By examining the spatial distribution of the eigenvalues, we observe that the relevant components are spread across nearly the entire image (See right panel of Fig. 12). In Figure (13) the results obtained for the stellar dataset are displayed: the eigenvalues associated to the input layer are plotted (black vertical lines) for a model trained with 100 epochs and a regularization coefficient for the eigenvalues set to $5\times 10^{-4}$.The test loss at the end of the training is essentially the same as that obtained for the case reported in Sec. (3.4) (not shown). However, the number of non-zero eigenvalues in the first layer is larger, although most of the peaks still lie in spectral regions known to be crucial for metallicity prediction. In other words the network detected another equivalent solution that makes use of a larger set of input components.
Figure 12: Distribution of the first layer eigenvalues for the MNIST dataset for a model trained with a smaller number of epochs (left) and spatial distribution of the eigenvalues on the input space (right).
12).
Figure 13: First layer eigenvalues obtained with fewer training epochs and a smaller regularization coefficient (black vertical lines), shown alongside the indices in the stellar spectra considered relevant for the metallicity prediction task (grey shadows)
We also evaluated the effect of different settings for other hyperparameters, such as the number of layers, the use of batch normalization, and the application of dropout. We did not observe any significant changes in the results under these variations. We would like to recall that our method is input-independent as it does not rely on the activation of the nodes to estimate their relevance (unlike methods such as LRP, for example). Additionally, the eigenvalues of the input layer are parameters directly tied to the input components and are unaffected by operations applied in the subsequent layers of the network. For these reasons, our method demonstrates strong robustness to architectural changes and to the use of techniques like batch normalization or dropout. Therefore, the hyperparameters were manually optimized from a restricted set of possible candidates by assessing the model’s performance on a validation set ($10$ of the training set).
Acknowledgements
This work is supported by #NEXTGENERATIONEU (NGEU) and funded by the Ministry of University and Research (MUR), National Recovery and Resilience Plan (NRRP), project MNESYS (PE0000006) "A Multiscale integrated approach to the study of the nervous system in health and disease" (DR. 1553 11.10.2022). We thank Francesco Belfiore and Michele Ginolfi for discussion about the astrophysical implications of our findings.